From the data shown below in the table about supply of alarm clocks, calculate the price elasticity from different points. Classify the elasticity of each point as elastic, inelastic and unit elastic.
Point Price Quantity Supplied Elasticity
A $8 50
B 9 70
C 10 80
D 11 88
C 12 95
E 13 100
1
Expert's answer
2019-03-08T07:11:06-0500

From point A to point B, price rises from $8 to $9, and quantity rises from 50 to 70. So:


% change in quantity= (70-50)/((70+50)÷2)×100=20/60×100=33.33


% change in price= (9-8)/((9+8)÷2)×100=1/8.5×100=11.76


Elasticity of Supply=(33.33%)/(11.76%)=2.83

The supply curve is elastic in this area; because its elasticity value is greater than one.

From point С to point D, the price rises from $10 to $11, and quantity rises from 80 to 88. So:


% change in quantity= (88-80)/((88+80)÷2)×100=8/84×100=9.52


% change in price= (11-10)/((11+10)÷2)×100=1/10.5×100=9.52


Elasticity of Supply=(9.52%)/(9.52%)=1


The supply curve has unitary elasticity in this area.

From point C to point E, the price rises from $12 to $13, and quantity rises from 95 to 100. So:


% change in quantity= (100-95)/((100+95)÷2)×100=5/97.5×100=5.13


% change in price= (13-12)/((13+12)÷2)×100=1/12.5×100=8.0


Elasticity of Supply=(5.13%)/(8.0%)=0.64


The supply curve is inelastic in this area.